OpenStudy (anonymous):

evaluate the integral from 0 to pie/8. (2x+sec2xtan2x)dx

OpenStudy (anonymous):

look in the book and see what the antiderivative of \[\sec^2(x)\tan^2(x)\] is

OpenStudy (anonymous):

oh wait sorry. this is \[\sec(2x)\tan(2x)\] yes?

OpenStudy (anonymous):

sec 2x=sec2x+tan2x

OpenStudy (anonymous):

tan2x=sec^22x

OpenStudy (anonymous):

:(\

OpenStudy (amistre64):

sec^2 tan^2 = (sec tan)^2

OpenStudy (anonymous):

no no easy way the derivative of secant is secant tangent

OpenStudy (anonymous):

so anti-derivative of secant tangent is secant.

OpenStudy (amistre64):

good notation does help ;)

OpenStudy (anonymous):

your anti-derivative is just \[x^2+\frac{1}{2}\sec(2x)\]

OpenStudy (anonymous):

the one half because you have \[\sec(2x)\] not \[\sec(x)\]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so now just plug in \[\frac{\pi}{8}\] and 0 and be happy